OpenStudy (anonymous):

Matt bought an oven for $30.oo . This price was 1/4 off the list price. What was the list price????????? :)

OpenStudy (anonymous):

if x=list price; (1/4)x=30 so x=120.

OpenStudy (anonymous):

nah. not that

OpenStudy (anonymous):

haha ok well im confused

OpenStudy (anonymous):

it's (1/4)x-x=30

OpenStudy (anonymous):

Hi :)

OpenStudy (anonymous):

hey JAke!

OpenStudy (anonymous):

helpppp ugh lol

OpenStudy (anonymous):

x-(1/4)x=30 (3/4)x=30 x=40

OpenStudy (anonymous):

oh ok that makes more sense lol thanks josiahh

OpenStudy (anonymous):

10 off 40 is 30 and 10 is 1/4 of 40

OpenStudy (anonymous):

can i ask another question???

OpenStudy (anonymous):

The perimeter of a rectangle is twice the sum of its length and width. The perimeter is 22 meters and its length is 2 meters more than twice its width. WHAT IS ITS LENGHT???

OpenStudy (anonymous):

these type of questions are so DAMN confusing to me :(

OpenStudy (anonymous):

HARTNN HELP ME :) lol

hartnn (hartnn):

(1/4) x =30 x= 120

hartnn (hartnn):

oh wait

OpenStudy (anonymous):

no, no second chances here... :)

hartnn (hartnn):

:P 1/4 th off the list price .... meand 1/4 th of total is discount ?

OpenStudy (anonymous):

lol are you sure lol

OpenStudy (anonymous):

no i dont need help with the first question i need help with the one i posted a few comments up lol

hartnn (hartnn):

ohh..perimeter?

OpenStudy (anonymous):

YUP lol

hartnn (hartnn):

that is easier

OpenStudy (anonymous):

The perimeter of a rectangle is twice the sum of its length and width. The perimeter is 22 meters and its length is 2 meters more than twice its width. WHAT IS ITS LENGHT??? p=2*(w+l) l=2+2w 32=2w+2l 2=-2w+l l=34

OpenStudy (anonymous):

no

OpenStudy (anonymous):

l=34/3

OpenStudy (anonymous):

no what lol

OpenStudy (anonymous):

the answer should be lenght=34/3

OpenStudy (anonymous):

2 more than twice width @hatnn

hartnn (hartnn):

length = 2 m more than 2width so l = 2w+2 22= 2(l+w) = 2(2w+2+w) = 6w+4 11=3w+2 3w= 9 w= 3 l=3+2 = 5

hartnn (hartnn):

now?

OpenStudy (anonymous):

oh snap the perimeter is 22 not 32 haha.

hartnn (hartnn):

sorry. l=2+2w = 2+ 2*3 = 2+6 = 8

hartnn (hartnn):

so length = 8

OpenStudy (anonymous):

oh ok awesome thanks guys you are lifesavers to me right now OMG :):):)

OpenStudy (anonymous):

hey so does that answer make sense then or is it not 100% correct yet lol

hartnn (hartnn):

length = 8 meters for sure

OpenStudy (anonymous):

AWESOME :)

OpenStudy (anonymous):

nice that was totally correct i just got an A+ YEAH FOR OPEN STUDY HELPERS FOR THE RETARDED BLONDES LOL

hartnn (hartnn):

i too am retarded, but not blonde :P

OpenStudy (anonymous):

lol i definitely am in you have not yet noticed haha

hartnn (hartnn):

oh, i did ...

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

darnit liz... you still playing that blonde card?!?!?!?